OpenStudy (anonymous):

9. The final method of solving systems of equations covered is called “multiplication,” but is actually a form of elimination that requires an extra step, multiplication. Given the following system of equations, explain how you would solve using elimination and give your final answer. a. x – y = 5 b. 2x + 6y = -14

OpenStudy (anonymous):

Multiply the entire top equation by -2 and add it to the bottom, so: x-y=5 -> -2x+2y=10 +2x+6y=-14 -------------------------- The X's cancel, so we have 8y=-4. Now solve for y first.